Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability

Expert replies
by sacx » Thu Jan 22, 2009 2:36 am
Darcy, Gina, Ray and Susan will be the only participants at a meeting. There will be three soft chairs in the room where the meeting will be held, and one hard chair. Darcy and Ray will arrive simultaneously, but Gina and Susan will arrive individually. The probability that Gina will arrive first is 1/3, and the probability that Susan will arrive first is 1/3. The probability that Gina will arrive last is 1/3, and the probability that Susan will arrive last is 1/3. Upon arriving at the meeting, each of the participants will select a soft chair, if one is available. If Darcy and Ray arrive and see only one unoccupied soft chair, they will flip a fair coin to determine who will sit in that chair. By what percent is the probability that Darcy will sit in a soft chair greater than the probability that Gina will sit in a soft chair

A. 50%
B 25%
C. 16 2/3%
D. 12.5%
E. 0%
SACX
Join the discussion
Source: — Problem Solving |

by Ian Stewart » Thu Jan 22, 2009 3:20 am
Considering that they're describing a fairly simple situation, there are far too many words in the question! Anyway, I posted a solution here:

www.beatthegmat.com/advanced-probabilit ... 13122.html
For online GMAT math tutoring, or to buy my higher-level Quant books and problem sets, contact me at ianstewartgmat at gmail.com

ianstewartgmat.com
Join the discussion

by gaggleofgirls » Thu Jan 22, 2009 8:56 am
To answer, we need to determine, the % that Gina gets a soft chair and the % that Darcy gets a soft chair.

If Gina arrives 1st or middle, she gets a soft chair, if she arrives last, she does not. 2 of the 3 times she can arrive, she gets the soft chair, so it is 2/3.

Darcy is very similar, if he arrives first, he (and Ray) arrive first, Darcy gets a soft chair, if he (and Ray) arrive middle, Darcy gets a soft chair, the only difference is that if he arrives last, he does not automatically get the hard chair like Gine, instead he has a 50% chance of getting the soft chair. 50% of 1/3 = 1/6 so Darcy's chance of getting the soft chair is 1/3+1/3+1/6 = 5/6

Now, what we need to know if the % difference between 5/6 and 2/3. I like to use a common denominator for this so it is the % diff between 5/6 and 4/6, which is what % greater is 5 than 4, so 25%.

Answer is B

-Carrie
Join the discussion

by krisraam » Thu Jan 22, 2009 12:10 pm
Answer id B -25%

My Approach is how many times does darcy or Gina get to sit on a Hard Chair.

First take Darcy

Darcy will sit in the Hard Chair only when Gina and Susan are first ans second( vice versa ) and Darcy has to lose the toss.

So the probabilty is :: 1/3(Gina Coming first) * 1/3( Susan coming second 1- (1/3(coming last) + 1/3(Coming first)) ) * 1/2 ( losing the toss to Ray) = 1/9 * 1/2

The same thing happens when Susan comes first

So 2* 1/9 * 1/2 = 1/9

The probablity Darcy gets to sit on a soft chair = 8/9

Now take Gina

Gina will sit on Hard chair only when Gina is the last person = 1/3

The probablity Gina gets to sit on softchair = 2/3

So % difference is (8/9-2/3)/(8/9) = 1/4 = 25%






Darcy, Gina, Ray and Susan will be the only participants at a meeting. There will be three soft chairs in the room where the meeting will be held, and one hard chair. Darcy and Ray will arrive simultaneously, but Gina and Susan will arrive individually. The probability that Gina will arrive first is 1/3, and the probability that Susan will arrive first is 1/3. The probability that Gina will arrive last is 1/3, and the probability that Susan will arrive last is 1/3. Upon arriving at the meeting, each of the participants will select a soft chair, if one is available. If Darcy and Ray arrive and see only one unoccupied soft chair, they will flip a fair coin to determine who will sit in that chair. By what percent is the probability that Darcy will sit in a soft chair greater than the probability that Gina will sit in a soft chair

A. 50%
B 25%
C. 16 2/3%
D. 12.5%
E. 0%
Join the discussion